Operations Portal is an EGI service provided by CCIN2P3,
co-funded by EGI.eu and EGI-Engage


Release Version : 3.3 - December 29th

For any contact, use this section: contact us


The Operations Portal is a central portal for the EGI operations management that offers a different capabilities, such as the broadcast tool, VO management facilities and various of dashboards (Security, VO and Operations) to facilitate infrastructure oversight.

Latest news

Dear VO members,

This is to inform you of the decommission process of computing services from IBERGRID. The following services will be fully decommissioned:

cream01-tic.ciemat.es (CREAM-CE)
apel01-tic.ciemat.es (gLite-APEL)


The timeline of this process:

Timeline
====
From 4th of May 2016
- The CE queues grid services will be definitively disabled.

4th of June 2016
- Resource hardware and computing services may become inaccessible without further notice to VO managers, users, etc. from this date.
- End of log retention period.


If you have any dough's or any issues please we will try to help you.

Best regards,
Antonio Juan Rubio Montero
Dear VO users and managers, the WMS/LB and top-BDII services hosted at eu-india-02.pd.infn.it and eu-india-03.pd.infn.it will be decommissioned by the end of May 2016.
Dear All,

We would like to announce that ICEAGE-CATANIA (NGI_IT) site has started the decommission procedure, which will follow this schedule :

29/04/2016 it will start queue draining 
30/04/2016 downtime will begin

It should be possible to recover data from the site until May 16th.

All the services sites will be following the same deadline. 
Best regards,
Emidio Giorgio - ICEAGE-CATANIA
Dear VO members,

This is to inform you of the decommission process of LIP-Coimbra site from IBERGRID. The following services will be fully decommissioned:

ui2.gridc.lip.pt (UI)		
ui1.gridc.lip.pt (UI)	
ce01.gridc.lip.pt (APEL)	
grid003.gridc.lip.pt (Site-BDII)	
ce01.gridc.lip.pt (CREAM-CE)	
storm.gridc.lip.pt (SRM)	
grid005.gridc.lip.pt (gLite-APEL)	
ce01.gridc.lip.pt (gLExec)	
ps01.gridc.lip.pt (net.perfSONAR.Latency)	
ps02.gridc.lip.pt (net.perfSONAR.Bandwidth)	
argus.gridc.lip.pt (emi.ARGUS)


The timeline of this process:

Timeline
====
From 27th of April 2016
- The all of its grid services will begin scheduled downtime until 1st of June, during which time VO managers and users can retrieve data.

2th of June 2016
- *** HARD DEADLINE for VO Managers to retrieve all data from storage elements. ***

3th of June 2016
- The resource centre status will be marked "suspended".
- Resource centre hardware and services may become inaccessible without further notice to VO managers, users, etc. from this date.

3th of August 2016
- End of log retention period.
- Resource centre will be marked "closed".

If you have any dough's or any issues please we will try to help you.

Best regards,
João Pina
Dear VO members,

This is to inform you of the decommission process of  LIP-Lisbon site from IBERGRID. The following services will be fully decommissioned:

ce02.lip.pt (computing element)	
se05.lip.pt  (storage) 
srm01.lip.pt (storage)


The timeline of this process:

Timeline
====
From 27th of April 2016
- The all of its grid services will begin scheduled downtime until 1st of June, during which time VO managers and users can retrieve data.

2th of June 2016
- *** HARD DEADLINE for VO Managers to retrieve all data from storage elements. ***

3th of June 2016
- The resource centre status will be marked "suspended".
- Resource centre hardware and services may become inaccessible without further notice to VO managers, users, etc. from this date.

3th of August 2016
- End of log retention period.
- Resource centre will be marked "closed".

If you have any dough's or any issues please we will try to help you.

Best regards,
João Pina
This is to notify you of the decommissioning of the following services hosted at UKI-SOUTHGRID-BRIS-HEP:

lcgce03.phy.bris.ac.uk.ac.uk (CREAM-CE)
lcgce04.phy.bris.ac.uk.ac.uk (CREAM-CE)

The following timeline will be followed for the decommissioning steps:

26/4/2016 - Announcement and EGI Broadcast
26/4/2016 - long+medium queues set to draining
28/4/2016 - Services will be place into downtime
29/4/2016 - Services will be decommissioned and removed from the GOCDB.

>>> More news <<<

Dear All,

This is to notify you of the decommissioning of the following services hosted at UKI-SCOTGRID-GLASGOW:

svr022.gla.scotgrid.ac.uk (WMS)
svr023.gla.scotgrid.ac.uk (WMS)
svr024.gla.scotfgrid.ac.uk (L&B)

The following timeline will be followed for the decommissioning steps:

22/4/2016 - Announcement and EGI Broadcast
9/5/2016 - Services will be place into downtime and be set to drain, downtime will be set to last 2 weeks.
23/5/2016 - Service will be decommissioned and removed from the GOCDB.

The process can be tracked via the following GGUS ticket:
https://ggus.eu/?mode=ticket_info&ticket_id=120973

Yours Sincerely,

Gareth
Dear site managers,

you get this message because your site supports at least one of the following VOs: 

geohazards.terradue.com
hydrology.terradue.com

For the following VOMS server: 

voms.ba.infn.it

a new certificate from a *new CA* has been installed, so all the services based on VOMS authentication need to be reconfigured with the new parameters available here:  

https://voms.ba.infn.it:8443/voms/geohazards.terradue.com/configuration/configuration.action
https://voms.ba.infn.it:8443/voms/hydrology.terradue.com/configuration/configuration.action

In particular the corresponding .lsc file should look like this: 

~# cat /etc/grid-security/vomsdir/geohazards.terradue.com/voms.ba.infn.it.lsc 
/DC=org/DC=terena/DC=tcs/C=IT/ST=Rome/L=Frascati/O=Istituto Nazionale di Fisica Nucleare/CN=voms.ba.infn.it
/C=NL/ST=Noord-Holland/L=Amsterdam/O=TERENA/CN=TERENA eScience SSL CA 3

~# cat /etc/grid-security/vomsdir/hydrology.terradue.com/voms.ba.infn.it.lsc  
/DC=org/DC=terena/DC=tcs/C=IT/ST=Rome/L=Frascati/O=Istituto Nazionale di Fisica Nucleare/CN=voms.ba.infn.it
/C=NL/ST=Noord-Holland/L=Amsterdam/O=TERENA/CN=TERENA eScience SSL CA 3

Please update your configuration as soon as possible. Reference is: 

https://wiki.egi.eu/wiki/HOWTO16_How_to_enable_a_Virtual_Organisation_on_a_EGI_Federated_Cloud

but feel free to ask for help to EGI Operations through GGUS if needed.

EGI Operations
We have experienced a series of network outages overnight, affecting the Tier 1 and other RAL based services. A member of the network team is on-site. There is no time to fix yet.
Dear all,

this ticket is to inform you that the AEGIS09-FTN-KM Resource Centre is going to be decommissioned. The downtime of all services will begin at 22nd April 2016 and will last till 3rd June 2016. All the services (CREAM-CE, SE, UI, WN) will be decommissioned on 15th of June 2016.

According to https://wiki.egi.eu/wiki/PROC11, we will use this (parent) ticket to trace all the required steps.

Best regards,

Petar Milic
Affected ROC/NGI: NGI_AEGIS
Affected Site: AEGIS09-FTN-KM
We will remove Castor storage disk pool GenScratch from use at site RAL-LCG2 (This is the RAL Tier1 site).

The timescale for this is:
Monday 23rd May 2016: Write access to the disk pool will be stopped. (This is just over 6 weeks from now).
Monday 20th June 2016: All access to the disk pool will be stopped. (This is 4 weeks after the stopping of write access).
Dear VO managers and users,

after many years of loyal service we will decommission desdemona (CreamCE with associated cluster) and cleopatra (BDII, APEL).
They are already in downtime (after a irreparable hardware failure) and we will remove them from GOCDB tomorrow (April 8, 2016).

Ophelia, our dCache data service will stay in production. New computing resources will be put into production soon (via a new CreamCE).

Cheers, Ralph
Dear WLCG site admins in EGI,
the latest version of the GOCDB has introduced WLCG-reserved scope tags:

wlcg, atlas, alice, lhcb, cms, tier1, tier2

For each site an initial set of relevant tags was determined and then applied
recursively to _all_ services provided by the site.

Please check the services of your site and _remove_ any spurious tags.

In particular, if your site has a Storage Element that is dedicated to a
subset of the experiments that you support, please ensure that the tags
for that service only list the proper subset.

Please open a ticket for the GOCDB if you think there is a mistake in the
set of tags applied to your _site_ (you cannot change them yourself).

Release notes:

https://wiki.egi.eu/wiki/GOCDB/notifications#v5.6_Released_For_Testing
The problem reported at the RAL Tier1 has been fixed. Services are now available. The site has been put in a warning in the GOC DB until 10:00 UTC tomorrow (30th March).
The RAL Tier1 is currently unavailable owing to a network problem. The problem started at around 19:00 UTC today (29th March). Staff are investigating. We do not yet have a time to fix.